OpenStudy (maddiie):

(-3,___) is a solution of y=4-3x Find the missing coordinate. -> I think that the answer is -5 but i'm confused. Can someone help?

OpenStudy (anonymous):

plug in -3 to the x part of the equation so y=2-3(-3) y=2-(-9) y=2+9 (subtracting a negative number makes a positive) y=11 so the solution is (-3,11)

OpenStudy (anonymous):

y=4-3(-3)=4+9=13

OpenStudy (anonymous):

answer is not -5!

OpenStudy (anonymous):

answer would have been -5 if x-coordinate was 3

OpenStudy (maddiie):

Okay i understand the way you've expained it it. Thank you

OpenStudy (anonymous):

click on good answer:)

OpenStudy (maddiie):

Where is that button

OpenStudy (anonymous):

its on the right of each post heading..

OpenStudy (maddiie):

Well i don't know if i'm dumb but i'm really blind. I can't see it. Give me the steps like a smal child plz... ;0)

OpenStudy (anonymous):

k...wat do u see above each post..there is a grey post heading...on that grey heading the poster's name and pic is on the left and the good answer button with a small medal pic is on the right..for some weird reason the text color in which "good answer" is written is white..but thers a link on the right side of the grey bar

OpenStudy (maddiie):

You know what that's not what i see on my end. What i see is the picture and name on the left and then the medal and shield looking thing right next to that. On my right it says 0 medals for this answer. Not nothing about good answer.

OpenStudy (anonymous):

yes after "0 medals for this answerer" hover ur mouse on that little area u will find a link...

OpenStudy (anonymous):

oh sorry I thought that it said 2- not 4-. sorry about that

OpenStudy (maddiie):

Nada

OpenStudy (maddiie):

Let's just give up. I suck dude... lol

OpenStudy (anonymous):

see attached picture...

OpenStudy (anonymous):

did u see?

OpenStudy (anonymous):

its now in blue color appearing...

OpenStudy (maddiie):

Funny on your screen i see the button but it doesn't show on my end there is no such button. Should i log off and log back on?

OpenStudy (anonymous):

yeps try..

OpenStudy (anonymous):

it will appear on the other's post..for u r screen it will appear on my post!

OpenStudy (maddiie):

i got it now.I sign off and came back

OpenStudy (maddiie):

lol.. geez

OpenStudy (anonymous):

k..good keep giving these for every user who satisfies u!! i also give u a medal since u have none..and took so pain to give me one...love u :)

OpenStudy (maddiie):

Thanks. Anything for someone who helps me with this Algrebra that i'm so bad at

OpenStudy (maddiie):

One medal for me. Yay!